(10%) Problem 6: Two lenses are mounted d= 29 cm apart on an optical bench. The focal length of the first lens is fı = 7.9 cmThe type of lens is not given.
Answer #1

Both lenses are positive so both focal lengths are convex lens. • - for lens } lens formula? / - - 1 -29 t = = £ 1 V 7.9 = v=10.86cm a u au -> u=-/29 - 10.86) cm = -18.14 6m E for lens @ Last - I v -18.14 - 5.6 5.6 V = between 8.1 cm. object & distan